# Towards Group-aware Search Success

### v1

@@ Abstract.
copyright: acmcopyrightjournalyear: 2022doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conferences title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYbooktitle: Woodstock ’18: ACM Symposium on Neural Gaze Detection, June 03–05, 2018, Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06copyright: none
@@ 3.1. Group-aware Search Success within Query
(4) $p(\mathbf{S}^{\text{GA}}|q)$ $=\prod_{g}{p(s|q,g)},$ (5) $=\prod_{g}{\Bigg{(}\sum_{t}^{\tau}{p(t|q,g)\cdot p(s|t,q)}\Bigg{)}}.$
@@ 3.2. Group-aware Search Success across Queries
    (6) $p(\mathbf{S}^{\text{GA}}_{\sum\prod})$ $=\sum_{q}p(\mathbf{S},q),$ (7) $=\sum_{q}p(\mathbf{S}|q)\cdot p(q),\quad\text{$q$ has no overlap}$ (8) $=\sum_{q}\bigg{(}\prod_{g}p(s|q,g)\bigg{)}\cdot p(q),$ (9) $=\sum_{q}\bigg{(}\prod_{g}\Big{(}\sum_{t}{p(t|q,g)\cdot p(s|t,q,g)}\Big{)}\bigg{)}\cdot p(q),$ (10) $=\sum_{q}\bigg{(}\prod_{g}\Big{(}\sum_{t}{p(t|q,g)\cdot p(s|t,q)}\Big{)}\bigg{)}\cdot p(q).\quad\text{if }s\perp\!\!\!\perp g\;|\;t,q$
@@
    (11) $p(\mathbf{S}^{\text{GA}}_{\prod\sum})$ $=\prod_{g}p(s|g),$ (12) $=\prod_{g}\bigg{(}\sum_{q}p(s,q|g)\bigg{)},$ (13) $=\prod_{g}\bigg{(}\sum_{q}p(s|q,g)\cdot p(q|g)\bigg{)},$ (14) $=\prod_{g}\bigg{(}\sum_{q}\Big{(}\sum_{t}{p(t|q,g)\cdot p(s|t,q,g)}\Big{)}\cdot p(q|g)\bigg{)},$ (15) $=\prod_{g}\bigg{(}\sum_{q}\Big{(}\sum_{t}{p(t|q,g)\cdot p(s|t,q)}\Big{)}\cdot p(q|g)\bigg{)}.$

### v2

@@ Abstract.
Search Success; Diversity; Quality-of-Service; Stochastic Ranking
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 2024 ACM SIGIR International Conference on the Theory of Information Retrieval; July 13, 2024; Washington DC, DC, USAbooktitle: Proceedings of the 2024 ACM SIGIR International Conference on the Theory of Information Retrieval (ICTIR ’24), July 13, 2024, Washington DC, DC, USAdoi: 10.1145/3664190.3672526isbn: 979-8-4007-0681-3/24/07ccs: Information systems Query intentccs: Information systems Information retrieval diversityccs: Information systems Relevance assessmentccs: Information systems Presentation of retrieval results
@@ 3.1. Group-aware Search Success within Query
(4) $p(\mathbf{S}^{\text{GA}}|q)$ $=\prod_{g}{p(s|q,g)},$
(5) $=\prod_{g}{\Bigg{(}\sum_{t}^{\tau}{p(t|q,g)\cdot p(s|t,q)}\Bigg{)}}.$
@@ 3.2. Group-aware Search Success across Queries
    (6) $p(\mathbf{S}^{\text{GA}}_{\sum\prod})$ $=\sum_{q}p(\mathbf{S},q),$
    (7) $=\sum_{q}p(\mathbf{S}|q)\cdot p(q),\quad\text{$q$ has no overlap}$
    (8) $=\sum_{q}\bigg{(}\prod_{g}p(s|q,g)\bigg{)}\cdot p(q),$
    (9) $=\sum_{q}\bigg{(}\prod_{g}\Big{(}\sum_{t}{p(t|q,g)\cdot p(s|t,q,g)}\Big{)}\bigg{)}\cdot p(q),$
    (10) $=\sum_{q}\bigg{(}\prod_{g}\Big{(}\sum_{t}{p(t|q,g)\cdot p(s|t,q)}\Big{)}\bigg{)}\cdot p(q).\quad\text{if }s\perp\!\!\!\perp g\;|\;t,q$
@@
    (11) $p(\mathbf{S}^{\text{GA}}_{\prod\sum})$ $=\prod_{g}p(s|g),$
    (12) $=\prod_{g}\bigg{(}\sum_{q}p(s,q|g)\bigg{)},$
    (13) $=\prod_{g}\bigg{(}\sum_{q}p(s|q,g)\cdot p(q|g)\bigg{)},$
    (14) $=\prod_{g}\bigg{(}\sum_{q}\Big{(}\sum_{t}{p(t|q,g)\cdot p(s|t,q,g)}\Big{)}\cdot p(q|g)\bigg{)},$
    (15) $=\prod_{g}\bigg{(}\sum_{q}\Big{(}\sum_{t}{p(t|q,g)\cdot p(s|t,q)}\Big{)}\cdot p(q|g)\bigg{)}.$